DIRECT SPECTRAL EVIDENCE OF SINGLE-AXIS ROTATION AND ORTHO-HYDROGEN-ASSISTED NUCLEAR SPIN CONVERSION OF CH_{3}F IN SOLID PARA-HYDROGEN

YUAN-PERN LEE, Department of Applied Chemistry and Institute of Molecular Science, National Chiao Tung University, Hsinchu 30010, Taiwan; YU-JONG WU, National Synchrotron Radiation Research Center, Hsinchu 30076, Taiwan; \underline{JON T. HOUGEN}, Optical Technology Division, National Institute of Standards and Technology, Gaithersburg, MD 20899-8441.

Numerous studies have shown that methane and its various deuterated isotopologs CH_{4−n}D_{n} (n = 0 − 4) can rotate relatively freely in solid parahydrogen. Less is known, however, of whether methyl groups can also carry out rotation (or internal rotation) in p-H_{2}, and even less is known of whether other large-amplitude motions involving hydrogen atoms (e.g., inversion or H transfer) can occur easily. The present talk, which is part of an attempt to systematically investigate these latter questions in National Chiao Tung University, presents infrared spectra for the nondegenerate (ν_{1}, ν_{2}, ν_{3}) and doubly degenerate (ν_{4}, ν_{5}, ν_{6}) vibrational fundamentals of CH_{3}F, together with a theoretical analysis strongly suggesting that CH_{3}F does indeed rotate about its symmetry axis (a axis) in p-H_{2}, but does not rotate about axes perpendicular to the symmetry axis. The theoretical analysis makes use of the fact that a-axis rotation is expected to give rise: (i) to energy levels of the form E = AK^{2}, (ii) to first-order Coriolis interactions in degenerate vibrational states (E states), and (iii) to different nuclear spin functions for rovibrational A (K = 0) and E (K = 1 and 2) states. Experimental values for the relatively slow E → A conversion rate in p-H_{2} have also been determined. The results for CH_{3}F will be compared with our earlier results for CH_{3}OH.